OpenStudy (anonymous):

Given A =1/2(b1 + b2)h, solve for h?

OpenStudy (anonymous):

I know that you have to isolate h.

OpenStudy (anonymous):

okay good now remember PEMDAS?

OpenStudy (anonymous):

yes sort of but would i distribute 1/2?

OpenStudy (anonymous):

yes, because you have to take away the P in PEMDAS first which is parentesis.

OpenStudy (anonymous):

ok so where would "A" go. Would you have to subtract it by 1/2

OpenStudy (anonymous):

its easier to show you then explain..will you let me?

OpenStudy (anonymous):

sure please do

OpenStudy (anonymous):

okay. A =1/2(b1 + b2)h A =(b1 + b2)h (2) 2*A= (b1+b2)h 2*A= (b1+b2)h /(b1+b2) \[2A \div (b1+b2)= h\]

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

ok so wait how did you get this again: A =(b1 + b2)h (2). did you distribute?

OpenStudy (anonymous):

I times two on both sides to get rid of 1/2

OpenStudy (anonymous):

to get 2A

OpenStudy (anonymous):

oh ok got it. I understand now thanks! :)

OpenStudy (anonymous):

Welcome! I hope that helped! ^^